6: An empty aerosol-spray can at room temperature (20 ◦C) is thrown into an incinerator
where the temperature reaches 605◦C. If the
gas inside the empty container was initially at
a pressure of 1 atm, what pressure did it reach
inside the incinerator? Assume the gas was at
constant volume and the can did not explode.
Answer in units of atm.
9: To what temperature must a sample of helium
gas be cooled from 111◦C to reduce its volume
from 4.7 L to 0.77 L at constant pressure?
Answer in units of K.
11: A 757 mL gas sample at STP is compressed to
a volume of 275 mL, and the temperature is
increased to 34◦C. What is the new pressure
of the gas in Pa?
Answer in units of Pa.
1
Expert's answer
2020-04-19T15:30:30-0400

6. p1/T1 = p2/T2

T1 = 20 + 273 = 293 K

T2 = 605 + 273 = 878 K

1 atm / 293 K = p2 / 878 K

p2 = 1 atm x 878 K / 293 K = 3 atm

9. V1 / T1 = V2 / T2

T1 = 111 + 273 = 384 K

4.7 L / 384 K = 0.77 L / T2

T2 = 384 K x 0.77 L / 4.7 L = 63 K

11. (p1 x V1) / T1 = (p2 x V2) / T2

V1 = 757 mL = 0.757 L; T1 = 273 K; p1 = 105 Pa

V2 = 275 mL = 0.275 L; T2 = 34 + 273 = 307 K

(105 Pa x 0.757 L) / 273 K = (p2 x 0.275 L) / 307 K

p2 = (105 Pa x 0.757 L x 307 K) / (273 K x 0.275 L) = 3 x 105 Pa